An aqueous KNO3 solution is made using 67.00 g of KNO3 diluted to a total solution volume of 1.60 L . Assume a density of 1.05 g/mL for the solution. What is the molarity?
Answer #1

67.00 g of KNO3 = mass / molar mass = 67.00 / 101.1 = 0.6627 mole.

volume of solution is 1.60 L .

molarity = 0.6627 * 1.0 / 1.60 = 0.414 M (answer)

density is needed during calculation of molality but not for molarity.